OpenStudy (anonymous):

The selling price, s, of an item is s = c + mc, where c is the cost of the item and m is the percent markup based on cost. What is the formula solved for m?

OpenStudy (anonymous):

\[ s = c+mc = 1c+mc = (1+m)c \]

OpenStudy (anonymous):

This is how you could get rid of \(c\).

OpenStudy (midhun.madhu1987):

Subtract c from both sides. Then Divide both sides by c and you get the value of m.

OpenStudy (anonymous):

When you have \(m+1=s/c\), you are close.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

@midhun.madhu1987 may you help me please

OpenStudy (midhun.madhu1987):